OpenStudy (tyrinzoebrown):

I need help, medals givin 3x+10>=4

OpenStudy (john_es):

\[3x+10\geq4\Rightarrow 3x\geq4-10\Rightarrow 3x\geq-6\Rightarrow x\geq-6/3\Rightarrow x\geq-3\]

OpenStudy (tyrinzoebrown):

Thanx john :)